# Distributed MIMO Networks with Rotary ULAs for Indoor Scenarios under Rician Fading

### v1

@@ Abstract
The Fifth-Generation (5G) wireless communications networks introduced native support for Machine-Type Communications (MTC) use cases. Nevertheless, current 5G networks cannot fully meet the very stringent requirements regarding latency, reliability, and number of connected devices of most MTC use cases. Industry and academia have been working on the evolution from 5G to Sixth Generation (6G) networks. One of the main novelties is adopting Distributed Multiple-Input Multiple-Output (D-MIMO) networks. However, most works studying D-MIMO consider antenna arrays with no movement capabilities, even though some recent works have shown that this could bring substantial performance improvements. In this work, we propose the utilization of Access Points (APs) equipped with Rotary Uniform Linear Arrays (RULAs) for this purpose. Considering a spatially correlated Rician fading model, the optimal angular position of the RULAs is jointly computed by the central processing unit using particle swarm optimization as a function of the position of the active devices. Considering the impact of imperfect positioning estimates, our numerical results show that the RULAs’s optimal rotation brings substantial performance gains in terms of mean per-user spectral efficiency. The improvement grows with the strength of the line-of-sight components of the channel vectors. Given the total number of antenna elements, we study the trade-off between the number of APs and the number of antenna elements per AP, revealing an optimal number of APs for the cases of APs equipped with static ULAs and RULAs.
@@
6G, Distributed MIMO, Location-Based Beamforming, Particle Swarm Optimization, Machine-Type Communications.
@@ I Introduction
The vast majority of works investigating the performance of D-MIMO networks consider fixed antenna arrays, i.e., antenna arrays with no movement capabilities. Nevertheless, the idea of antenna arrays that can move has gained attention among the research community since some recent works have shown that antenna movements can substantially improve the quality of wireless links [7, 8, 9, 10, 11, 12, 13], as discussed in the following subsection.
@@ I-B Contributions and Organization of the Paper
  • We propose a scheme for joint configuration of the angular positions of all the RULAs in the system to maximize the mean per-user achievable Spectral Efficiency (SE). The scheme relies on known estimates the positions of the MTDs and Particle Swarm Optimization (PSO).
@@
  • We propose a positioning error model and define different levels of positioning accuracy based on the requirements specified by the Third Generation Partnership Project (3GPP) standards. The system’s performance is evaluated as a function of the positioning accuracy. Our numerical results show that the optimal rotation of the RULAs brings performance gains even when the system’s positioning accuracy is poor.
@@
Table I lists the acronyms used throughout this paper alphabetically. This paper is organized as follows. Section II presents the system model. Section III introduces the positioning error model. Section IV describes the proposed mechanism for optimizing the angular position of the RULAs. Section V presents and discusses the numerical results. Finally, Section VI concludes the paper.
@@ II System Model
We consider a square coverage area with dimensions $l\times l\;\text{m}^{2}$. $Q$ APs serve the coverage area, each equipped with a RULA of $S$ half-wavelength spaced antenna elements. The APs are at height $h_{\text{AP}}$. A large number $K_{\text{total}}$ of single-antenna MTDs are distributed on the coverage area. The positions of the MTDs are fixed for a relatively large period. From this large number of MTDs, a random subset of $K$ MTDs are active and seek to transmit data in the uplink in each time slot. Let $(x_{k},y_{k})$ denote the coordinates of the $k$-th MTD. For simplicity, we consider that the antenna elements of all MTDs are positioned at the same height $h_{\text{MTD}}$ [4, 5]. The system model is illustrated in Fig. 1.
We consider fully centralized processing: all the APs are connected to a common CPU through fronthaul connections. The CPU is responsible for performing the following signal-processing tasks:
@@
    Computation of the estimates of the positions of the $K_{\text{total}}$ MTDs using an indoor localization method.
@@
    MTD’s activity detection and identification.
@@
    Computation of the optimal angular positions of the RULAs in each time slot based on the locations of the active MTDs.
@@
    Channel State Information (CSI) acquisition through pilot sequences.
@@
    Computation of centralized receive combining vectors based on the channel estimates.
@@
Further details on each of these tasks will be provided in the subsequent sections of this work.
@@ II-A Channel Model
The spatial covariance matrices can be (approximately) modeled using the Gaussian local scattering model [20, Sec. 2.6]. Specifically, the $s$-th row, $m$-th column element of the correlation matrix is
@@
Figure 2: Illustration of the three-phase RA scheme for 6 active MTDs [21].
II-B Three-Phase Random Access
Herein, we assume that the CPU has estimates of the positions of all MTDs in the network. We adopt a variation of the three-phase Random Access (RA) scheme introduced in [21] for scheduled massive access. Each of its phases is described below:
  1. 1.
    Phase 1: The active users transmit non-orthogonal uplink pilots for user identification and activity detection.
  2. 2.
    Phase 2: The CPU identifies the set of active users and determines the scheduling for the uplink transmissions. Based on the positions of the $K$ devices that will be active in each time slot, the CPU computes the optimal angular positions for the RULAs for each time slot. Finally, the system broadcasts a common downlink feedback message to assign each user a time slot and an orthogonal pilot.
  3. 3.
    Phase 3: The MTDs simultaneously transmit their orthogonal pilots and data in their scheduled slot. As the CPU jointly controls the APs, the RULAs perform their rotations in each time slot.
The three-phase RA scheme is illustrated in Fig. 2. Phase 1 corresponds to Slot 1, Phase 2 is performed during Slots 1 and 2, and Phase 3 comprises Slot 2 onwards. In Slot 1, 6 MTDs become active and transmit signals $\textbf{s}_{k},\;k\in\{1,\ldots,6\}$ intended for activity detection and user identification. Then, $K=2$ MTDs are scheduled for each subsequent time slot. Each of them is assigned one of the $K=2$ orthogonal pilot sequences, $\{\bm{\phi}_{1},\bm{\phi}_{2}\}$, and transmit their data signal, $\textbf{x}_{k},\;k\in\{1,\ldots,6\}$.
In this work, we consider that the scheduling of MTDs for Phase 3 is random. Moreover, we assume MTC scenarios where the devices are stationary, such as smart utility meters [22] or video surveillance cameras [23]. The location of an MTD is obtained as soon as it is registered to the network [^2] Indoor positioning technologies are discussed in [24, 25, 26]..
We assume that the duration of a time slot is long enough to accommodate the computation of the optimal angular positions of the RULAs, their rotations, and the uplink transmissions from the $K$ MTDs. The orthogonal pilots transmitted by the MTDs are used for CSI estimation. Then, the estimated channel coefficients are used for data decoding, as described in the next subsection.
II-C Signal Model
Herein, we consider Phase 3 of the RA scheme: an uplink scenario where $K$ active MTDs simultaneously transmit signals to the $Q$ APs. Let $M=QS$ denote the total number of receive antenna elements. The collective vector of wireless channel coefficients between the $k$-th MTD and the $Q$ APs is $\textbf{h}_{k}=[\textbf{h}_{k1}^{T},\textbf{h}_{k2}^{T},\ldots,\textbf{h}_{kQ}^{T}]^{T}\in\mathbb{C}^{M\times 1}$.
@@ II-C Signal Model
The receive combining matrix V is computed as a function of the matrix of estimated channel vectors $\hat{\textbf{H}}\in\mathbb{C}^{M\times K}$, $\hat{\textbf{H}}=[\hat{\textbf{h}}_{1},\ldots,\hat{\textbf{h}}_{K}]$. In this work, we adopt the centralized Zero Forcing (ZF) combining scheme [^3] Centralized ZF for distributed massive MIMO in indoor industrial scenarios was also adopted in [27]. Note that the SE performance obtained with ZF approaches that with MMSE in the high SINR regime [28].. The receive combining matrix is computed as [29]
@@ II-D Performance Metrics
II-E Imperfect CSI Model
The estimated channel vector of the $k$-th MTD, $\hat{\textbf{h}}_{k}\in\mathbb{C}^{M\times 1}$, can be modeled as the sum of the true channel vector plus a random error vector as [30, 31]
@@ II-E Imperfect CSI Model
where $\tilde{\textbf{h}}_{k}\sim\mathcal{CN}(\textbf{0},\sigma_{\text{csi}}^{2}\textbf{I}_{M})$ is the vector of channel estimation errors. Note that the true channel realizations and the channel estimation errors are uncorrelated.
The parameter $\sigma_{\text{csi}}^{2}$ indicates the quality of the channel estimates. Let $\rho=p/\sigma^{2}_{n}$ denote the transmit SNR. Assuming orthogonal pilot sequences during the uplink data transmission phase and least squares channel estimation, it can be modeled as a decreasing function of $\rho$ as [32]
$\sigma_{\text{csi}}^{2}=\dfrac{1}{K\rho}.$ (15)
III Positioning Error Model
As mentioned in Section II-B, the CPU utilizes the information about the positions of the MTDs to compute the optimal rotations of the RULAs. However, the estimates of the positions are not perfect in practical systems. In this section, we propose a mathematical model for the positioning error and relate it to the positioning accuracy requirements proposed by 3GPP.
Considering that all the MTDs are positioned at the same height $h_{\text{MTD}}$, the imperfect positioning impairment refers to the uncertainty on the position of the MTDs only on the $(x,y)$ axes. Let $\textbf{p}_{k}=(x_{k},y_{k})$ denote the true position of the $k$-th MTD, and $\hat{\textbf{p}}_{k}=(\hat{x}_{k},\hat{y}_{k})$ denote the estimated position. The positioning error vector associated to the $k$-th MTD becomes
@@ III Positioning Error Model
where $x_{e,k}=x_{k}-\hat{x}_{k}$ and $y_{e,k}=y_{k}-\hat{y}_{k}$ are the $x$ and $y$ components of the positioning error vector, respectively. We can rewrite the positioning error vector in polar form as
@@
where $r_{e,k}=\sqrt{x_{e,k}^{2}+y_{e,k}^{2}}$ and $\theta_{e,k}=\tan^{-1}(y_{e,k}/x_{e,k})$ are the magnitude and phase of the positioning error vector, respectively.
We assume the positioning error follows a bivariate Gaussian distribution with mean $\bm{\mu}=[\textbf{0}\;\textbf{0}]^{T}$ and covariance matrix $\bm{\Sigma}=\sigma_{e}^{2}\textbf{I}_{2}$ [33]. Thus, $x$ and $y$ components of the positioning error vector follow a Normal distribution:
@@
Moreover, the magnitude of the positioning error vector follows a Rayleigh distribution:
@@
where $\sigma_{e}^{2}$ is the scaling parameter. Finally, the angle of the positioning error vector follows a uniform distribution:
@@
Let $r_{e,k}$, i.e., the radius of the positioning error vector, represent the positioning accuracy. Considering the bivariate Gaussian model, the positioning accuracy follows the Rayleigh distribution. Thus, we have
@@
In 3GPP standards, the positioning accuracy requirements are typically specified in terms of the 95% quantile [34]. In 3GPP Rel-16, the positioning accuracy requirement specified for indoor scenarios was $3$ m [35]. In 3GPP Rel-17, this requirement was further reduced to $20$ cm [36]. The positioning accuracy requirement for beyond-5G and 6G networks in indoor factory scenarios is expected to be in the order of a few centimeters [37].
@@ IV Optimal Angular Positions of the RULAs
In each time slot, $K$ distinct MTDs are active. As shown in our previous work [17], a distinct optimal angular position exists for the RULA for each subset of $K$ positions of active MTDs. Moreover, high-speed industrial servo motors can change their positions in a few milliseconds [38, 39], thus they can track slow-to-moderate network dynamics in practice.
@@
The optimal rotations of the RULAs $\theta_{q}\,;\forall q\in\{1,\ldots,Q\},$ are jointly computed by the CPU as a function of $\hat{\textbf{p}}_{k},\;\forall k\in\{1,\ldots,K\}$. Owing to the large number of variables to be jointly optimized and the high non-linearity of this problem, it is solved using particle swarm optimization (PSO), which will be presented in the next subsection.

### v2

@@ Abstract
The Fifth-Generation (5G) wireless communications networks introduced native support for Machine-Type Communications (MTC) use cases. Nevertheless, current 5G networks cannot fully meet the very stringent requirements regarding latency, reliability, and number of connected devices of most MTC use cases. Industry and academia have been working on the evolution from 5G to Sixth Generation (6G) networks. One of the main novelties is adopting Distributed Multiple-Input Multiple-Output (D-MIMO) networks. However, most works studying D-MIMO consider antenna arrays with no movement capabilities, even though some recent works have shown that this could bring substantial performance improvements. In this work, we propose the utilization of Access Points (APs) equipped with Rotary Uniform Linear Arrays (RULAs) for this purpose. Considering a spatially correlated Rician fading model, the optimal angular position of the RULAs is jointly computed by the central processing unit using particle swarm optimization as a function of the location of the active devices. Considering the impact of imperfect location estimates, our numerical results show that the RULAs’s optimal rotation brings substantial performance gains in terms of mean per-user spectral efficiency. The improvement grows with the strength of the line-of-sight components of the channel vectors. Given the total number of antenna elements, we study the trade-off between the number of APs and the number of antenna elements per AP, revealing an optimal number of APs for the cases of APs equipped with static ULAs and RULAs.
@@
6G, Distributed MIMO, Location-Based Beamforming, Machine-Type Communications, Particle Swarm Optimization.
@@ I Introduction
The vast majority of works investigating the performance of D-MIMO networks consider fixed antenna arrays, i.e., antenna arrays with no movement capabilities. Nevertheless, the idea of antenna arrays that can move has gained attention among the research community [7, 8, 9, 10, 11, 12, 13]. According to recent works [12, 14, 13], moving and/or rotating antennas can take advantage of spatially varying channel conditions within a confined area. Indeed, by altering the position and/or orientation of the antenna, better channel conditions can be achieved. In MTC scenarios, devices are typically deployed in fixed locations, while the surrounding environment may change over time, leading to slowly varying wireless channels. In use cases without mobility, narrowband MTC offers limited time and frequency diversity, which restricts the potential for improving data rates and transmission reliability. In such situations, antennas with movement capabilities are a promising solution for achieving higher spatial diversity gains. The related works will be discussed in more detail in the following subsection.
@@ I-B Contributions and Organization of the Paper
  • We propose a scheme for joint configuration of the angular positions of all the RULAs in the system to maximize the mean per-user achievable Spectral Efficiency (SE). We assume that the network obtains estimates of the locations of the active devices (i.e., the devices that seek to transmit data in the uplink) before the data transmission phase and using some indoor localization method. The estimates are utilized to compute the optimal angular positions of the RULAs of the APs using the Particle Swarm Optimization (PSO).
@@
  • We propose a localization error model and define different levels of localization accuracy based on the requirements specified by the Third Generation Partnership Project (3GPP) standards. The system’s performance is evaluated as a function of the localization accuracy. Our numerical results show that the optimal rotation of the RULAs brings performance gains even when the system’s localization accuracy is poor.
@@
Table I lists the acronyms used throughout this paper alphabetically. This paper is organized as follows. Section II presents the system model. Section III introduces the localization error model. Section IV describes the proposed mechanism for optimizing the angular position of the RULAs. Section V presents and discusses the numerical results. Finally, Section VI concludes the paper.
@@ II System Model
Similar to simulation setups of many relevant and related works (see, for instance, [4, 19, 20, 21]), we consider a square coverage area with dimensions $l\times l\;\text{m}^{2}$. $Q$ APs serve the coverage area, each with a RULA of $S$ half-wavelength spaced antenna elements. The APs are at height $h_{\text{AP}}$. The $Q$ APs jointly serve $K$ MTDs that are active simultaneously and transmit data in the uplink. The locations of the MTDs are fixed, that is, we consider a scenario with no mobility. Let $(x_{k},y_{k})$ denote the coordinates of the $k$-th MTD. For simplicity, we consider that the antenna elements of all MTDs are located at the same height $h_{\text{MTD}}$ [4, 5]. The system model is illustrated in Fig. 1.
We consider fully centralized [^2] In this work, for the sake of simplicity and to highlight the novelty of our proposed framework, we assume that all APs are perfectly calibrated and synchronized, a common assumption in several relevant works (e.g., [19, 4, 20, 21]). Note that imperfect synchronization can affect any D-MIMO system in both cases of APs equipped with static or rotary/movable antennas. Nevertheless, there is no indication that any alternatives would be more or less sensitive to synchronization. processing: all the APs are connected to a common CPU through high-capacity fronthaul connections. The CPU is responsible for performing the following signal-processing tasks:
@@
    MTD’s activity detection and identification.
@@
    Computation of the estimates of the locations of the active MTDs using an indoor localization method.
@@
    Computation of the optimal angular positions of the RULAs based on the estimated locations of the active MTDs.
@@
    Computation of the Channel State Information (CSI) of the links between any AP and any MTD using pilot sequences.
@@
    Computation of centralized receive combining vectors based on the global CSI knowledge.
@@
Further details on each of these tasks are provided in the subsequent sections of this work.
@@ II-A Channel Model
The spatial covariance matrices can be (approximately) modeled using the Gaussian local scattering model [24, Sec. 2.6]. Specifically, the $s$-th row, $m$-th column element of the correlation matrix is
@@
II-B Signal Model
Herein, we consider an uplink scenario where $K$ active MTDs simultaneously transmit signals to the $Q$ APs. Let $M=QS$ denote the total number of receive antenna elements. The collective vector of wireless channel coefficients between the $k$-th MTD and the $Q$ APs is $\textbf{h}_{k}=[\textbf{h}_{k1}^{T},\textbf{h}_{k2}^{T},\ldots,\textbf{h}_{kQ}^{T}]^{T}\in\mathbb{C}^{M\times 1}$.
@@ II-B Signal Model
The receive combining matrix V is computed as a function of the matrix of estimated channel vectors $\hat{\textbf{H}}\in\mathbb{C}^{M\times K}$, $\hat{\textbf{H}}=[\hat{\textbf{h}}_{1},\ldots,\hat{\textbf{h}}_{K}]$. In this work, we adopt the centralized Zero Forcing (ZF) combining scheme [^3] Centralized ZF for distributed massive MIMO in indoor industrial scenarios was also adopted in [25]. Note that the performance obtained with ZF approaches that with MMSE in the high SINR regime [26], with the advantage of not requiring statistical knowledge of the noise variance or interference.. The receive combining matrix is computed as [27]
@@ II-C Performance Metrics
II-D Imperfect CSI Model
Assuming $\tau_{p}$ orthogonal pilot sequences during the uplink data transmission phase, such that $\tau_{p}\geq K$, and least squares channel estimation, the true channel realizations and the channel estimation errors are uncorrelated [28]. Then, the estimated channel vector of the $k$-th MTD, $\hat{\textbf{h}}_{k}\in\mathbb{C}^{M\times 1}$, can be modeled as the sum of the true channel vector plus a random error vector as [29, 30, 28]
@@ II-D Imperfect CSI Model
where $\tilde{\textbf{h}}_{k}\sim\mathcal{CN}(\textbf{0},\sigma_{\text{csi}}^{2}\textbf{I}_{M})$ is the vector of channel estimation errors.
The parameter $\sigma_{\text{csi}}^{2}$ indicates the quality of the channel estimates. Let $\rho=p/\sigma^{2}_{n}$ denote the transmit SNR. We also assume that the duration of the uplink pilot transmission phase is equal to $\tau_{p}$ symbols. Then, the variance of the channel estimation errors can be modeled as a decreasing function of $\rho$ as [29, 30, 28]
$\sigma_{\text{csi}}^{2}=\dfrac{1}{\tau_{p}\rho}.$ (15)
Note that the channel estimation error depends only on the uplink transmit power, average receive noise power and number of orthogonal pilots, thus it is the same for all devices. The consequence of imperfect CSI is a slight decrease in the mean per-user achievable SEs compared to the case of perfect CSI for both the cases of static ULAs and RULAs.
Figure 2: Illustration of the proposed framework for the optimization of the angular positions of the RULAs and uplink data transmission. Note that the phases have different duration.
II-E Proposed Framework
In this subsection, we describe our proposed framework for the optimization of the rotation of the $Q$ APs and subsequent uplink data transmission. We assume that the same subset of devices is active during a period of time that spans multiple consecutive time slots. For instance, the devices could be video surveillance cameras [31] or smart utility meters [32]. Therefore, we assume that the optimization of the angular positions of the RULAs is performed before the data transmission phase, thus the angular positions of the RULAs is kept constant during multiple consecutive time slots. Inspired by the three-phase scheduled random access scheme from [33], our framework has the following phases:
  1. 1.
    Active MTDs, i.e., MTDs seeking to send data to the network, transmit non-orthogonal uplink pilots for activity detection.
  2. 2.
    The network identifies the set of active MTDs and, utilizing some indoor localization techniques [34, 35, 36], obtains estimates of the locations of the active MTDs.
  3. 3.
    The network assumes pure LoS propagation and utilizes the estimated locations of the MTDs and location-based beamforming to compute its optimal rotation and/or position. The RULAs have their angular positions updated accordingly.
  4. 4.
    The network broadcasts a common downlink feedback message to assign each device an orthogonal pilot sequence.
  5. 5.
    The MTDs transmit their orthogonal pilot sequences and data during multiple time slots. The uplink orthogonal pilots are used to compute the CSI estimates shown in (14) for each coherence time interval, which are then used to compute the ZF receive combining vectors in (11).
The proposed framework is illustrated in Fig. 2. Note that phase 5 extends over $T$ time slots. Within each time slot, the $K$ active devices transmit simultaneously. Each time slot is as long as the channel’s coherence time interval and contains $\tau_{c}$ symbols. The first $\tau_{p}$ symbols of each time slot are used for the transmission of the orthogonal pilot sequences, while the remaining $\tau_{d}=\tau_{c}-\tau_{p}$ symbols are utilized for uplink data transmission. The numerical results presented in this work, that is, the mean per-user achievable SE presented in Section II-II-C, correspond to the performance achieved during phase 5.
III Localization Error Model
As mentioned in Section II-E, the CPU utilizes the information about the locations of the MTDs to compute the optimal rotations of the RULAs. However, the estimates of the locations are not perfect in practical systems. In this section, we propose a mathematical model for the localization error and relate it to the localization accuracy requirements proposed by 3GPP.
Considering that all the MTDs are located at the same height $h_{\text{MTD}}$, the imperfect localization impairment refers to the uncertainty on the location of the MTDs only on the horizontal plane. Let $\textbf{p}_{k}=(x_{k},y_{k})$ denote the true location of the $k$-th MTD, and $\hat{\textbf{p}}_{k}=(\hat{x}_{k},\hat{y}_{k})$ denote the estimated location. The localization error vector associated to the $k$-th MTD becomes
@@ III Localization Error Model
where $x_{e,k}=x_{k}-\hat{x}_{k}$ and $y_{e,k}=y_{k}-\hat{y}_{k}$ are the $x$ and $y$ components of the localization error vector, respectively. We can rewrite the localization error vector in polar form as
@@
where $r_{e,k}=\sqrt{x_{e,k}^{2}+y_{e,k}^{2}}$ and $\theta_{e,k}=\tan^{-1}(y_{e,k}/x_{e,k})$ are the magnitude and phase of the localization error vector, respectively.
Aiming at generality (that is, to not introduce any additional assumption or bias related to indoor localization), we model the localization error as a bivariate Gaussian distribution since it is the least informative [^4] In information theory and statistics, “least informative” refers to the distribution with the maximum entropy (i.e., uncertainty) for a given variance. distribution for any given variance [37]. Specifically, we assume the localization error follows a bivariate Gaussian distribution with mean $\bm{\mu}=[\textbf{0}\;\textbf{0}]^{T}$ and covariance matrix $\bm{\Sigma}=\sigma_{e}^{2}\textbf{I}_{2}$ as in [38]. Thus, $x$ and $y$ components of the localization error vector follow a Normal distribution:
@@
Moreover, the magnitude of the localization error vector follows a Rayleigh distribution:
@@
where $\sigma_{e}^{2}$ is the scaling parameter. Finally, the angle of the localization error vector follows a uniform distribution:
@@
Let $r_{e,k}$, i.e., the radius of the localization error vector, represent the localization accuracy. Considering the bivariate Gaussian model, the localization accuracy follows the Rayleigh distribution. Thus, we have
@@
In 3GPP standards, the localization accuracy requirements are typically specified in terms of the 95%-quantile [39]. The $F$-quantile of a Rayleigh distribution with scaling parameter $\sigma_{e}$ is given by
$Q(F;\sigma_{e})=\sigma_{e}\sqrt{-2\ln(1-F)}.$ (23)
Considering the proposed localization error model and based on the 3GPP specifications, we list examples of localization accuracy requirements for pessimistic, reasonable, and optimistic scenarios and their corresponding values of $\sigma_{e}^{2}$ in Table II. In 3GPP Rel-16, the localization accuracy requirement specified for indoor scenarios was $3$ m [40]. In 3GPP Rel-17, this requirement was further reduced to $20$ cm [41]. The localization accuracy requirement for beyond-5G and 6G networks in indoor factory scenarios is expected to be in the order of a few centimeters [42].
TABLE II: Examples of localization accuracy requirements for pessimistic, optimistic, and reasonable scenarios [40, 41, 42].
$\phantom{-}\sigma_{e}^{2}$ 95% Quantile Scenario
[…]
$-7.8$ dB $1$ m Reasonable
$-28$ dB $10$ cm Optimistic
$-45$ dB $1$ cm Optimistic
@@ IV Optimal Angular Positions of the RULAs
In each time slot, $K$ distinct MTDs are active. Our previous work [18] shows a distinct optimal angular position for the RULA for each subset of $K$ locations of active MTDs. Moreover, high-speed industrial servo motors can change their angular positions within a few milliseconds [^5] In this work, we assume that the rotation of the RULAs is performed only once and before the uplink data transmission phase. Consequently, the time required to rotate the RULAs does not yield channel aging or any other impact on the uplink data transmission. Nevertheless, timing requirements are crucial for the operation of our proposed system, and we intend to study them more thoroughly in future extensions of this work. with very high precision [43, 44]. Thus, they can track slow-to-moderate network dynamics in practice. Moreover, assuming the ULA is implemented as a microstrip patch antenna array, similar to the prototypes studied in [45, 46, 47, 48], it will be lightweight. Hence, it will not impose a significant payload burden on the servo motor.
@@
The optimal rotations of the RULAs $\theta_{q}\,;\forall q\in\{1,\ldots,Q\},$ are jointly computed by the CPU as a function of $\hat{\textbf{p}}_{k},\;\forall k\in\{1,\ldots,K\}$. This computation can be written as the following optimization problem:
maximize $f(\theta_{1},\ldots,\theta_{Q}\;|\;\hat{\textbf{p}}_{k},\;k\in\{1,\ldots,K\})$ (25)
subject to $0\leq\theta_{1},\ldots,\theta_{Q}\leq\pi,$
where $f(\cdot)$ is the objective function to be maximized. Owing to the large number of variables to be jointly optimized and the high non-linearity of this problem, it is solved using particle swarm optimization (PSO), which will be presented in the next subsection.